if a+b+c=0, then the roots of the equation 4ax^2+3bx+2c=0, where a,b,c are real numbers are

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
Given that:a+b+c = 0
a^2/(a^2 – bc) + b^2/(b^2 – ca) + c^2/(c^2 – ba).............(1)
Substituting in the eqn. (1)
a = -b-c or a^2 = (b+c)^2
we get,
2(a^2 +b^2 + ab)/(a^2 +b^2 + ab)
=> 2
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)